# Transfer Learning for Spatial Autoregressive Models

### v1

@@ 4 Simulation
We generate the data according to Models (2.1) and (2.2). In these models, $n_{0}=256,n_{1}=n_{2}=\cdots=n_{K}=100,p=1,q=200,K=20$. The covariates $\mathbf{x}_{i}^{(k)}$ are generated by multivariate normal distribution with covariance matrix $\Sigma\in\mathbb{R}^{q\times q}$, whose the $(j,j^{\prime})$-th entry is denoted as $\Sigma_{j,j^{\prime}}$ for all $0\leq k\leq K$. We consider three designs for $\Sigma$: (1) $\Sigma_{j,j^{\prime}}=1$ for $j=j^{\prime}$ and $\Sigma_{j,j^{\prime}}=0$ otherwise, (2) $\Sigma_{j,j^{\prime}}=0.5^{|j-j^{\prime}|}$, (3) $\Sigma_{j,j^{\prime}}=0.9^{|j-j^{\prime}|}$. We consider two different distributions for the i.i.d. error terms: (1) the standard normal distribution $N(0,1)$, and (2) the $t_{2}$ distribution. The spatial units are assumed to be located on a square grid, and we set up several candidate spatial weight matrices as follows: (1) the first $W_{1}$ is the spatial weight matrix where spatial units interact with their neighbors on the left and right sides; (2) the second $W_{2}$ is the spatial weight matrix where spatial units interact with their neighbor above and below; (3) the rest of the spatial weight matrices are assumed to be matrices where spatial units interact with their second-nearest, third-nearest, and so on, denoted as $W_{3},\ldots,W_{N}$, where $N<\sqrt{n_{0}}$. In each dataset, we randomly draw spatial weight matrices from these candidate matrices. The weight matrices ${\{W_{k,l}\}}_{k\in\{0\}\cup[K],l\in[p]}$ are drawn randomly from $\{W_{1},\ldots,W_{N}\}$ without replacement. In Models (2.1) and (2.2), we consider coefficients settings as follows. In the target Model (2.1), $\lambda_{0}$ $=0.4$ and $\beta_{0}$ $=(\mathbf{1}_{3}^{\intercal},\mathbf{0}_{q-3}^{\intercal})^{\intercal}\in\mathbb{R}^{q}$. In Model (2.2),
@@
$\mbox{RMSE}=\sqrt{\frac{1}{R}\sum_{r=1}^{R}\|{\theta_{0}}-{\widehat{\theta}}^{[r]}\|^{2}},$ (4.1)
@@
Figure 1: RMSE of Three Types of Covariates Designs with Normal Error Distribution, and $H=0,5,10$
Note: The size of informative auxiliary datasets $|\mathcal{A}|$ increases from 00 to $20$.
Table 1: The RMSE of Various Estimators with Standard Normal Distribution Error Terms
@@
Notes: The RMSE of $\widehat{\lambda}$ is the RMSE of the estimated coefficients of the $\lambda_{0}$. The total RMSE is the RMSE of all the estimated coefficients of the covariates. “NA” indicates that “glmtrans” is not applicable for the estimation of $\lambda_{0}$.
5 Empirical Study
In recent years, amounts of Geographic Information Science (GIS) data have been used in election campaigns (Peters et al.,, 2004). Geography is the common denominator of where voters live, how residents engage in the voting process, how officials manage elections, and where elections campaign runs. Based on population data and geographic information, election district boundaries are resettled to draft multiple proposed plans and demonstrate the effects of several approaches, for instance, moving a border to divide a community of interest. Political campaigns make full use of geographic information and related tools to obtain underlying support from voters, especially the ones from swing states.
Shaping voter preference in swing states is an essential problem, especially for the prediction target. Including spatial information in data analysis brings a novel insight into the prediction. In our empirical study, we focus on the 2020 U.S. presidential election results of each county for swing states by incorporating the spatial relationship among counties. We try to predict the election result using the polls from the 2016 election U.S. presidential election and some other demographic data with geographic information. It is worth noting that swing states have undergone historical shifts over time. The process of identifying swing states in previous elections typically involves an assessment of the closeness of the vote margins in each state. This analysis incorporates historical election outcomes, opinion polls, political trends, recent developments since the previous election, and the specific attributes, strengths, or vulnerabilities of the candidates in contention. We select six swing states for consideration from mainstream journals and news media, including Florida, Georgia, Michigan, Minnesota, North Carolina, and Ohio.
The digital boundary definitions of the United States congressional district are available at https://cdmaps.polisci.ucla.edu/, and the cartographic boundary shape files of counties in the USA are available at the United States Census Bureau (https://www.census.gov/geographies/mapping-files/time-series/geo/cartographic-boundary.2020.html). The spatial weight matrix is constructed using above two geographic datasets. We apply the “queen” contiguity-based spatial weights and row-normalized. The “queen” criterion defines neighbors as spatial units sharing a common edge or a common vertex. The response is the difference of support at the county level for presidential candidates from both parties in 2016, available at https://github.com/tonmcg/US_County_Level_Election_Results_08-20. The county-level demographic information forms the predictors, collected at https://www.kaggle.com/datasets/benhamner/2016-us-election. Among all states, we exclude Alaska, Hawaii, and Washington, D.C.. After regular preprocessing, there are 3106 counties and 45 county-level predictors, among 48 U.S. states. Then we add the pairwise interaction terms between predictors as new predictors. Finally, there are 1035 county-level predictors.
@@ 5 Empirical Study
Each swing state is regarded as a target, and all the other states are considered as data sources. We run the proposed estimation method, TranSAR, using the polls from the 2016 U.S. presidential election and some other demographic data with geographic information. We also estimate the SAR model with Lasso penalty (due to high dimensional covariates) using only the target datasets (denoted as SAR) to compare the empirical performance. Due to the inherent stochastic nature of the resampling procedure for the algorithm, we conduct 20 replications for both methods. The average prediction RMSE of the county-level vote results of the county-level 2020 U.S. presidential election is compared at Table 2. For all swing states, the proposed TranSAR method performs better than the traditional SAR in terms of RMSE.
Table 2: The Average RMSE of the Predicted County-level Rate of Support in the 2020 Presidential Election.
@@
Then, we compare the predicted rate of state level support and the true one in $2020$ election, showing in Table 3. It demonstrates that the transfer learning under the SAR framework is able to effectively improve the state-level prediction.
Table 3: Predicted Bias of the Predicted State-level 2020 Election Results.
@@
The calculated ratios of correct predictions out of the twenty election result forecasts are collected in Table 4. As shown in Table 4, the method we proposed exhibits better performance, especially in predicting election outcomes in Michigan and Minnesota. Simultaneously, it is noteworthy that for the election predictions in Georgia, neither of the methods works well. It can be explained by the historical context: from $1972$ to $2016$, Georgia consistently supported the Republican Party, except for Democratic candidates from the southern region. However, the state has become an increasingly competitive battleground. In 2020, the Democratic candidate, Joe Biden, secured a victory over Donald Trump by a narrow margin of $0.2\%$. This marks a significant departure from the history, as the Democrats secured the presidential election with a slim advantage. Evidently, Georgia has become an outlier in this regard. It is evident that predicting the ultimate election outcome for a single state is significantly more challenging than forecasting the rate of state-level support. Nevertheless, we have made commendable progress in this regard. To provide a comprehensive demonstration of the superiority of the proposed method, we treat each state as the target once a time. Subsequently, we conduct both the tranSAR and SAR processes, and the outcomes have been summarized into Table LABEL:tab:_emp_compare_ele_result_for_all_states for reference, which can be found in LABEL:sec:_addtional_empirical_results. In conclusion, our tranSAR method consistently outperforms the traditional SAR method in the majority of cases, yielding superior prediction results.
@@
In addition, we attempt to visualize transferable states, conducting a statistical analysis of the results after $20$ repetitions. For a swing state, we record the frequency of occurrences for all transferable states and depict states with more than $10$ occurrences (50%) on the map in green color, see Figure 2.
@@
As observed in the figure, some transferable states exhibit geographical information signals. However, for some targets, the spatial or geographical correlations may not be readily apparent. This is a commonplace occurrence. With the development of the economy, culture, and scientific technology, regional interconnections are no longer confined solely to spatial proximity. Latent economic and political affiliations become pivotal factors. We have also created a total visualization map, which is provided in Figure LABEL:fig_emp_transferable_states_map_top_50_all_states at LABEL:sec:_addtional_empirical_results for reference. In this study, we treat each state in the United States as the target once a time, and employ transfer learning techniques similar to those described earlier to construct these maps.

### v2

@@ 4 Simulation
We generate the data according to Models (1) and (2). In these models, $n_{0}=256,n_{1}=n_{2}=\cdots=n_{K}=100,p=1,q=200,K=20$. The covariates $\mathbf{x}_{i}^{(k)}$ are generated by multivariate normal distribution with covariance matrix $\Sigma\in\mathbb{R}^{q\times q}$, whose the $(j,j^{\prime})$-th entry is denoted as $\Sigma_{j,j^{\prime}}$ for all $0\leq k\leq K$. We consider three designs for $\Sigma$: (1) $\Sigma_{j,j^{\prime}}=1$ for $j=j^{\prime}$ and $\Sigma_{j,j^{\prime}}=0$ otherwise, (2) $\Sigma_{j,j^{\prime}}=0.5^{|j-j^{\prime}|}$, (3) $\Sigma_{j,j^{\prime}}=0.9^{|j-j^{\prime}|}$. We consider two different distributions for the i.i.d. error terms: (1) the standard normal distribution $N(0,1)$, and (2) the $t_{2}$ distribution. The spatial units are assumed to be located on a square grid, and we set up several candidate spatial weight matrices as follows: (1) the first $W_{1}$ is the spatial weight matrix where spatial units interact with their neighbors on the left and right sides; (2) the second $W_{2}$ is the spatial weight matrix where spatial units interact with their neighbor above and below; (3) the rest of the spatial weight matrices are assumed to be matrices where spatial units interact with their second-nearest, third-nearest, and so on, denoted as $W_{3},\ldots,W_{N}$, where $N<\sqrt{n_{0}}$. In each dataset, we randomly draw spatial weight matrices from these candidate matrices. The weight matrices ${\{W_{k,l}\}}_{k\in\{0\}\cup[K],l\in[p]}$ are drawn randomly from $\{W_{1},\ldots,W_{N}\}$ without replacement. In Models (1) and (2), we consider coefficients settings as follows. In the target Model (1), $\lambda_{0}$ $=0.4$ and $\beta_{0}$ $=(\mathbf{1}_{3}^{\intercal},\mathbf{0}_{q-3}^{\intercal})^{\intercal}\in\mathbb{R}^{q}$. In Model (2),
@@
$\mbox{RMSE}=\sqrt{\frac{1}{R}\sum_{r=1}^{R}\|{\theta_{0}}-{\widehat{\theta}}^{[r]}\|^{2}},$ (15)
@@
Figure 2: RMSE of Three Types of Covariates Designs with Normal Error Distribution, and $H=0,5,10$.
@@
Table 1: The RMSE of Various Estimators with Normal Distribution Error Terms. The RMSE of $\widehat{\lambda}$ is the RMSE of the estimated coefficients of the $\lambda_{0}$. The total RMSE is the RMSE of all the estimated coefficients of the covariates. “NA” indicates that “glmtrans” is not applicable for the estimation of $\lambda_{0}$.
5 An Application to U.S. Presidential Election Prediction
In this section, we apply the proposed method in Algorithm 2 to empirical data from the U.S. presidential election. As shown in Section 2.1, the response variable $Y^{(k)}$ is the difference of support rates in the county level, and covariates $\mathbf{X}^{(k)}$ include demographic and geographica data, such as including total population, total housing units, poverty people population, etc. With the observed spatial weight matrix $\mathbf{W}^{(k)}$, each swing state is regarded as a target with $k=0$, and all the other states are considered as sources $k\in[47]$. The target model (1) and sources models (2) could formulated as
$Y^{(k)}=\sum_{l=1}^{p}\lambda_{l0}^{(k)}W_{l}^{(k)}Y^{(k)}+\sum_{j=1}^{q}X^{(k)}_{j}\beta^{(k)}_{j0}+V^{(k)},\ k\in\{0\}\cup[K].$
@@
We run the proposed estimation method, TranSAR, using the polls from the 2016 U.S. presidential election and some other demographic data with geographic information. We also estimate the SAR model with Lasso penalty (due to high dimensional covariates) using only the target datasets (denoted as SAR) to compare the empirical performance. Due to the inherent stochastic nature of the resampling procedure in the algorithm, we conduct multiple replications (20 times) for both methods to obtain a stable result. The average prediction RMSE of the county-level vote results of the county-level 2020 U.S. presidential election is compared at Table 2. For all swing states, the proposed TranSAR method performs better than the traditional SAR in terms of RMSE, which demonstrates the necessarity of transfer learning in the SAR framework.
@@
Table 2: Average Prediction RMSE of the Predicted County-level Rates of Support in the 2020 Presidential Election.
@@
Then, we compare the prediction rate of state-level support and the true one in $2020$ election, showing in Table 3. It demonstrates that the transfer learning under the SAR framework is able to effectively improve the state-level prediction.
@@
Table 3: Prediction Bias of the Predicted State-level 2020 Election Results.
The calculated ratios of correct predictions out of the 20 election result forecasts are collected in Table 4. As shown in Table 4, the method we proposed exhibits better performance, especially in predicting election outcomes in Michigan and Minnesota. Simultaneously, it is noteworthy that for the election predictions in Georgia, neither of the methods works well. It can be explained by the historical context: from $1972$ to $2016$, Georgia consistently supported the Republican Party, except for Democratic candidates from the southern region. However, the state has become an increasingly competitive battleground. In 2020, the Democratic candidate, Joe Biden, secured a victory over Donald Trump by a narrow margin of $0.2\%$. This marks a significant departure from the history, as the Democrats secured the presidential election with a slim advantage. Evidently, Georgia has become an outlier in this regard. It is evident that predicting the ultimate election outcome for a single state is significantly more challenging than forecasting the rate of state-level support. Nevertheless, we have made commendable progress in this regard.
@@
Next, we attempt to visualize transferable states, conducting a statistical analysis of the results after multiple repetitions. For a swing state, we record the frequency of occurrences for all transferable states and depict states with more than 50% occurrences on the map in green color. As observed in Figure 3, some transferable states exhibit geographical information signals. However, for some targets, the spatial or geographical correlations may not be readily apparent. This is a commonplace occurrence. With the development of the economy, culture, and scientific technology, regional interconnections are no longer confined solely to spatial proximity. Latent economic and political affiliations become pivotal factors.
@@
In addition, to provide a comprehensive demonstration of the superiority of the proposed TranSAR method, we treat each state as the target once a time. Subsequently, we conduct both the tranSAR and SAR procedures, and the outcomes have been summarized in Supplementary Materials. In conclusion, our tranSAR method consistently outperforms the traditional SAR method in the majority of cases, yielding superior prediction results.
Motivated by public interest, we also predict the outcome of the 2024 U.S. presidential election for all states using demographic and geographical data in 2022, based on the trained tranSAR model from the 2020 U.S. polling data along with other demographic and geographical data. We use the 2022 demographic data as covariates to predict county-level support rates. And we use the 2020 total vote counts of county-level as weights to calculate state-level support rates (because in the U.S., the “winner-takes-all” system weights each county by its vote count). We focus on the predictions of election outcomes for eight “swing states” following the rule based on whether there was a change in the supported party over the past three elections. Therefore, we select Arizona, Georgia, Florida, Pennsylvania, Michigan, Wisconsin, Ohio, and Iowa as swing states. For non-swing states, we use the results from their last three presidential elections as the outcome for this election, while for swing states, we use our tranSAR method for prediction. Considering the randomness inherent in the algorithm, we conduct multiple rounds of analysis using our TranSAR procedure. In these analyses, if the Democratic Party receive support in more than half of the iterations, the swing state was classified as supporting the Democratic Party; otherwise, it was classified as supporting the Republican Party. We obtain the supporting party for each state (see Figure 4) and calculate the final support vote counts combined with the electoral votes. The final prediction result shows that the Democratic Party receive 309 electoral votes, exceeding the threshold of 269 votes. Therefore, we predict that the outcome of the 2024 U.S. presidential election will favor the Democratic Party.
Figure 4: Predicted Party Support by States in 2024. The striped states are the swing states whose election outcomes predicted by our TranSAR method.
6 Theoretical results
In this section, we will establish an estimation convergence rate for the proposed $\mathcal{A}$-TranSAR estimator by assuming $\underline{n}\equiv\min_{k\in[K]\cup\{0\}}n_{k}\rightarrow\infty$. Subsection 6.1 is for general results and Subsection 6.2 is based on the 2SLS estimation method. Finally, we prove the consistency of the transferable source detection algorithm in Subsection 6.3.
For theoretical analysis, we define the population form for the above estimator as
$\omega^{\mathcal{A}}=\underset{\omega\in\mathbb{R}^{p+q}}{\arg\min}\sum_{k\in\mathcal{A}}\alpha_{k}\bar{\mathit{\ell}}_{1}^{(k)}(\omega),$ (16)
[…]
$\omega^{(k)}=\underset{\omega\in\mathbb{R}^{p+q}}{\arg\min}\ \bar{\mathit{\ell}}_{1}^{(k)}(\omega).$ (17)
6.1 Convergence rate when $\mathcal{A}_{h}$ is known
Assumption 1 (Network structure).
The spatial weight matrices $W_{l}^{(k)}(k\in\{0\}\cup[K],l\in[p])$ are non-stochastic with zero diagonals, and $p$ is fixed and does not depend on $n_{0}$ or $n_{k}$. $S^{(k)}(\lambda)=I_{n_{k}}-\sum_{l=1}^{p}\lambda_{l}^{(k)}W_{l}^{(k)}$ is invertible for all $\lambda\in\Lambda$, where the parameter space $\Lambda$ is compact. Matrices $W_{l}^{(k)}$ and $S^{(k)}(\lambda)^{-1}$ are uniformly bounded in both row and column sum norms for all $n_{k},k\in[K]\cup\{0\},l\in[p]$ and $\lambda\in\Lambda$. And for all large enough $n_{k}$, $\|W_{l}^{(k)}S^{(k)}(\lambda)^{-1}\|_{F}/\sqrt{n_{k}}$ is bounded away from zero and infinity uniformly over $k\in[K]\cup\{0\},l\in[p]$ and $\lambda\in\Lambda$.
  1. (1)
    Denote $\mathbf{X}^{(0)}\equiv{(\mathbf{x}_{1}^{(0)},\dots\mathbf{x}_{n_{0}}^{(0)})}^{\intercal}$ and the covariance matrix of $\mathbf{x}_{i}^{(0)}$ for each $i$ is $\Sigma_{0}$. For any ${\beta}\in\mathbb{R}^{q},{\mathbf{x}_{i}^{(0)}}^{\intercal}{\beta}\ (i\in[n_{0}])$, are $i.i.d.$ $\kappa_{0}\|{\beta}\|_{2}$ sub-Gaussian with mean zero, and $0<\underline{\psi_{0}}=\lambda_{\min}(\Sigma_{0})\leq\lambda_{\max}(\Sigma_{0})=\overline{\psi_{0}}<\infty$.
  2. (2)
    The error terms $V_{i}^{(0)}\ (i\in[n_{0}])$ are $i.i.d.$sub-Gaussian with mean zero and variance $\sigma^{2}$.
  3. (3)
    As $n_{0}$rightarrow +$\infty$+ ∞, $\mathbb{P}\left(\lambda_{\min}(\frac{(\widehat{\mathbb{X}}^{(0)})^{\intercal}\widehat{\mathbb{X}}^{(0)}}{n_{0}})\geq\phi_{0}\right)\rightarrow 1$, where $\phi_{0}$ is a positive constant.
Assumption 3 (Convergence rates of the first step estimator).
The first step estimator $\widehat{\omega}$ minimizing (4) satisfies $\|\widehat{\omega}-\omega^{\mathcal{A}}\|_{1}=O_{\mathbb{P}}{(a_{n_{\mathcal{A}}}^{(1)})}$, and $n_{\mathcal{A}}^{-1}\|{{\mathbb{X}}^{\mathcal{A}}}(\widehat{\omega}-\omega^{\mathcal{A}})\|_{2}^{2}=O_{\mathbb{P}}(a_{n_{\mathcal{A}}}^{(2)})$, where $a_{n_{\mathcal{A}}}^{(1)}\rightarrow 0$ and $a_{n_{\mathcal{A}}}^{(2)}\rightarrow 0$ as $n_{\mathcal{A}}\rightarrow\infty$, and ${\mathbb{X}}^{\mathcal{A}}$ is stack of $\mathbb{X}^{(k)}$ for $k\in\mathcal{A}$.
Assumption 4 (Instruments).
When $n_{0}$ is large enough, the IV matrix
$Q^{(0)}=\left(\mathbf{X}^{(0)},W_{1}^{(0)}\mathbf{X}^{(0)},\ldots,W_{q}^{(0)}\mathbf{X}^{(0)}\right)$ (18)
satisfies $0<\psi_{*}^{(0)}\leq\frac{1}{\sqrt{n_{0}}}\|Q^{(0)}\|_{2}\leq{\psi^{(0)}}^{*}<\infty$ for some constants $\psi_{*}^{(0)}$ and ${\psi^{(0)}}^{*}$. There exist constant coefficient vectors ${\pi}_{j}\in\mathbb{R}^{d}$ for $j\in[p+q]$ such that $\max_{j=1,\dots,p+q}\|\widehat{\pi}_{j}-{\pi}_{j}\|_{2}=O_{\mathbb{P}}(\sqrt{\log(d)/n_{0}})$ and $\|\pi_{j}\|_{2}^{2}$ is uniformly bounded over $j$ and $n_{0}$, where $\widehat{\pi}_{j}\ (j\in[p+q])$ are defined in Eq. (7).
Remark 4.
Theorem 1.
(Convergence rate of the $\mathcal{A}$-TranSAR estimator) Let the parameter space be $\Theta=\Lambda\times\left\{{\beta}\mid\|{\beta}\|_{2}<M\right\}$ for some positive constant $M$. Under Assumptions 1-4, if the loss function $\operatorname{\mathit{\ell}}_{1}$ is twice differentiable, $\sup_{k\in\mathcal{A}}\|{\Sigma^{\mathcal{A}}}^{-1}\Sigma^{(k)}\|_{1}=\check{C}<\infty$, $s\frac{\log{q}}{n_{0}}=o(1)$, $q\asymp d$, and $\lambda_{\delta}\gtrsim\sqrt{\log{(q)}/n_{0}}$, where
[…]
Denote $S=\operatorname{support}(\theta_{0})$ and $s=|S|$. Then
$\|\widehat{\theta}-\theta_{0}\|_{2}^{2}=O_{\mathbb{P}}\left((\sqrt{\frac{\log{q}}{n_{0}}}h)\wedge(s\frac{\log{q}}{n_{0}})\wedge h^{2}+a_{n_{\mathcal{A}}}^{(2)}\right).$ (21)
Remark 5.
6.2 Convergence rate under the 2SLS loss
In this section, we verify the requirements in the above theorem for the 2SLS estimation and simplify the convergence rate of the $\mathcal{A}$-TranSAR estimator. As in Eq. (8), let $\operatorname{\mathit{\ell}}_{1}(\cdot\mid\omega)$ be the loss function of the 2SLS. To elaborate, let $Y^{\mathcal{A}}$ denote the stack consisting of $Y^{(k)}$ for all $k\in\mathcal{A}$, i.e., $Y^{\mathcal{A}}={({Y^{(k)}}^{\intercal},k\in\mathcal{A})}^{\intercal}$. $\mathbf{X}^{\mathcal{A}}$ and ${\mathbb{X}}^{\mathcal{A}}$ are defined similarly. Then the objective function (4) could be rewritten as
$\frac{1}{2n_{\mathcal{A}}}\|Y^{\mathcal{A}}-\widehat{\mathbb{X}}^{\mathcal{A}}\omega\|_{2}^{2}+\lambda_{\omega}\|\omega\|_{1},$ (22)
where $\widehat{\mathbb{X}}^{(k)}$ is defined similarly as $\widehat{\mathbb{X}}^{(0)}$ in Subsection 3.1, and $\widehat{\mathbb{X}}^{\mathcal{A}}$ is their stack. Here the IV matrices are $Q^{(k)}$ and $Q^{\mathcal{A}}$. We impose several similar conditions on $\mathfrak{D}^{(k)}$ ($k\in\mathcal{A}$) in the following.
Assumption 5.
(Regular conditions for design matrices and error terms on the $k$-th source) For source $k\in\mathcal{A}$, we have the following assumptions: (i) For the design matrix $\mathbf{X}^{(k)}=(\mathbf{x}_{1}^{(k)},\dots\mathbf{x}_{n_{k}}^{(k)})^{\intercal}$ and any ${\beta}\in\mathbb{R}^{q},{\mathbf{x}_{i}^{(k)}}^{\intercal}{\beta}\ (i\in[n_{k}])$ are $i.i.d.\kappa_{k}\|{\beta}\|_{2}$ sub-Gaussian with mean zero for certain constant $\kappa_{k}$, where $\max_{k\in\mathcal{A}}\kappa_{k}<\infty$. (ii) The error terms $V_{i}^{(k)}\ (i\in[n_{k}])$ are $i.i.d.$ sub-Gaussian with mean zero and variance $\sigma_{k}^{2}$. (iii) As $n_{k}\rightarrow+\infty$, $\mathbb{P}\left(\lambda_{\min}\left(\frac{{{\widehat{\mathbb{X}}}^{(k)\intercal}}{\widehat{\mathbb{X}}}^{(k)}}{n_{k}}\right)\geq\phi_{k}\right)\rightarrow 1$, where $\phi_{k}$ is a positive constant.
Assumption 6 (Instruments for the $k$-th source).
For the instruments
$Q^{(k)}=\left(\mathbf{X}^{(k)},W_{1}^{(k)}\mathbf{X}^{(k)},\ldots,W_{q}^{(k)}\mathbf{X}^{(k)}\right),\ k\in\mathcal{A}\cup\{0\},$ (23)
when $n_{k}$, the number of rows of $Q^{(k)}$, is large enough, $0<\psi_{*}^{(k)}\leq\frac{1}{\sqrt{n_{k}}}\|Q^{(k)}\|_{2}\leq{\psi^{(k)}}^{*}<\infty$ for some constants $\psi_{*}^{(k)}$ and ${\psi^{(k)}}^{*}$. There exist constant coefficient vectors $\pi_{j}^{(k)}\in\mathbb{R}^{d}$ for $j\in[p+q]$ such that $\max_{j=1,\dots,p+q}\|\hat{\pi}_{j}-\pi_{j}^{(k)}\|_{2}=O_{\mathbb{P}}(\sqrt{\log(d)/n_{k}})$ and $\|\pi_{j}^{(k)}\|_{2}$ is uniformly bounded over $j$ and $k$.
Theorem 2.
$\|\widehat{\theta}-\theta_{0}\|_{2}^{2}=O_{\mathbb{P}}\left((\sqrt{\frac{\log{q}}{n_{0}}}h)\wedge({\frac{s\log{q}}{n_{0}}})\wedge h^{2}+(\frac{\log{n_{A}}}{n_{\mathcal{A}}}h^{2})\wedge\frac{\log^{3}{n_{A}}}{n_{\mathcal{A}}^{3}}+\frac{s\log{q}}{n_{\mathcal{A}}}\right).$ (24)
6.3 Detection consistency
Assumption 7 (Identifiability condition).
The tolerance level $h=o(1)$ as the minimal sample size $\underline{n}\equiv\min_{k\in[K]\cup\{0\}}n_{k}\rightarrow\infty$. For any $k\in[K]\cup\{0\}$ and $r\in[3]$, $q\log(q)\|\widetilde{\theta}^{(k)}-\widehat{\theta}^{(k,r)}\|_{1}=o_{\mathbb{P}}(1)$. For $k\in\mathcal{A}^{c}\equiv[K]-\mathcal{A}$, we control the signal power as
$\frac{\sigma^{2}}{2n_{0}}\operatorname{tr}\{[(S(\tilde{\lambda}^{(k)})-S)S^{-1}]^{\intercal}[(S(\tilde{\lambda}^{(k)})+S)S^{-1}]\},$ (25)
and the gap between $\theta_{0}$ and $\widetilde{\theta}^{(k)}$,
$\varphi_{k}\equiv\sqrt{\frac{1}{n_{0}}}(\|\theta_{0}-\widetilde{\theta}^{(k)}\|_{1}+1)^{2}=o(1),$ (26)
Remark 6.
Conditions (7) and (26) guarantee a suitable gap between the population-level coefficients in set $\mathcal{A}^{c}$ and the true coefficient of the target model. Intuitively, assuming an appropriate signal strength is necessary to ensure the consistency of estimators. On the one hand, a very weak signal makes it challenging for the algorithm to differentiate informative and uninformative sources. On the other hand, an excessively strong signal can lead the first step estimator to mimic the model structure of the source, thereby hindering accurate identification.
Theorem 3 (Detection consistency of $\widehat{\mathcal{A}}$).
If Assumptions (1)-(4) and (7) are satisfied, $\|\beta_{0}\|_{2}<\infty$ and $\frac{\log q}{n_{0}}=o(1)$, then the automatically detected transferable index set $\widehat{\mathcal{A}}$ in Eq. (14) is consistent, i.e.,
[…]